# Teaching Data Science Students to Sketch Privacy Designs through Heuristics (Extended Technical Report)

### v1

@@ Abstract
00footnotetext:  00footnotetext: Equal contributions.00footnotetext: This report is an extended version of a IEEE S&P 2025 paper. We include Table XIV to further prove the improved communication efficiency of heuristic-guided privacy design sketches (Section 5.3.3). This table was skipped in the original version due to space constraints.
@@ 1 Introduction
We conducted a detailed experiment to validate the effectiveness of the heuristic-based approach (Section 5). We recruited a total of 24 participants, with each participant creating a few sketches and interpreting the sketches from other participants. We found that participants who received heuristic-based instructions were able to cover 30% more of the privacy-related decisions in their sketches than the participants without exposure to these instructions. The final sketches from the participants who received heuristic-based instruction are also more readable, with an increase of 77% in interpretation accuracy.
@@ 2.2 Education for Privacy Design
Current privacy design education primarily focuses on privacy concepts [16] and design methodologies [17, 18], with most educators using case studies [17, 19] as the main form of instruction. For instance, some classes include a privacy review procedure [20, 19], where students assess smartphone privacy notices, identify issues, and propose design improvements in brief essays. Organizations such as the National Institute of Standards and Technology (NIST) [21] and the International Association of Privacy Professionals (IAPP) [22] also provide privacy design education by offering standards, certifications, and online learning resources, which provide practical guidance and structure for both academic and professional learning. Additionally, educators have expanded the instruction of privacy design to a broader audience, including industry professionals and policymakers [23]. In contrast, our work aims to complement existing curricula by exploring methods to teach data science students how to sketch privacy designs, equipping them with privacy design communication skills.
@@ 2.3 Diagrams for Privacy Design
Practitioners often use a combination of diagrams to communicate privacy considerations in system designs [3, 4, 43]. Research has shown that these diagrams help developers visualize system architecture by simplifying complex design elements, making it easier to communicate and design privacy-aware applications [4]. For example, when tasked with communicating the privacy design of an IoT application for diabetes treatment and monitoring, developers used Data flow diagrams (DFDs) [44] and Unified Modeling Language (UML) [45] to break down the flow of sensitive health data into manageable components, with annotations clarifying the encryption mechanism [4].
While existing studies shed light on the use of diagrams for communicating privacy considerations in system design, their focus has been on professional developers. However, how to teach these diagramming techniques to novice data science students remains unclear. Our work aims to address this gap by using sketches as the lens to investigate the specific challenges students face when communicating privacy design through diagrams and provide insights on targeted guidance to address these difficulties.
@@ 2.4 Developer Support for Privacy Design
Prior research has explored many approaches to support experienced developers in privacy design [46, 47]. For instance, some studies have offered insights into how to help developers make privacy design choices by identifying the common challenges in understanding and implementing privacy considerations [48, 49]. Additionally, various tools have been developed to support developers by embedding privacy practices directly into development workflows. For example, Coconut, an Android Studio plugin, helps developers create privacy-friendly apps by requiring privacy annotations and assisting them in organizing privacy-related details [50]. Similarly, PARROT helps developers create privacy-aware IoT applications by providing interactive privacy annotations and guidance [4].
@@ 4.1.1 Design
Second, we offered preset attribute options to help novice students explore the design space systematically (Figure 3 and 4). Instead of describing each DataAction in natural language, students could select predefined attribute options. For example, when specifying data source to Collect, students could choose ‘direct’ for newly generated data or ‘secondary’ for pre-existing data.
@@ 4.3.3 Observations
All participants completed the warm-up within 20 minutes (Mean = 18.5 minutes, SD = 1.2 minutes) and reached satisfaction with each design within 15 minutes (Mean = 13.6 minutes, SD = 2.9 minutes). First, we observed a significant reduction in completion time with the vocabulary-based approach (14.3 v.s. 18.6 minutes in Section 4.2.4), as participants reported that the warm-up session helped them become familiar with the method. Additionally, there was a clear preference for heuristics over vocabulary-based design. For example, participants who started with heuristic-guided sketching often intentionally applied the multi-layered heuristic when sketching vocabulary-based designs. However, we did not observe a tendency to re-apply visual vocabulary to heuristic-based tasks. T13 noted, “I felt freer to move beyond ellipses and boxes and to organize my ideas in a more trackable way.”
@@ 5.1 Study Overview
Study Procedure and Apparatus. We followed a between-subject study design. Participants received either vocabulary-based or heuristic-based guidance for their tasks. They completed tasks on an iPad with a stylus using a sketch board application (e.g., Notability [59]).
@@
Then, each participant was asked to interpret two sketches created by previous participants (Table 5.2.4). Finally, we interviewed participants about what hindered or assisted them in their creation and interpretation tasks.
@@
Annotating Visual Sketches. Comparing high-dimensional sketches is challenging. To address this, we developed a codebook (Appendix Table XI) to annotate privacy-related design decisions covered in the sketches. We then used the annotated decisions to assess the design coverage in sketch creation and the communication effectiveness between the sketch creators and interpreters (Sections 5.2, 5.3).
@@ 5.2.1 Data Collection
We presented participants with a list of privacy-related design decisions (Table XI) and asked them to describe each decision they had included, annotating the corresponding content on their sketches. We didn’t set time limits for this process. Participants could skip any decisions they felt were not covered but could not modify their original designs.
@@ 5.2.2 Data Analysis
There is no perfect sketch for the creation task, as participants may have different designs for the same scenario, which may evolve during sketching. So, we use the codebook (Table XI) to systematically capture the unique design decisions reflected in each sketch. To assess the quality of the sketches, we documented the privacy design decisions annotated by participants, counted the total number of decisions included in each sketch, and calculated the coverage percentage of each design decision across all sketches in each condition.
Next, we conducted a thematic analysis [70] of post-task interview transcripts and notes we took during study. First, two researchers independently coded eight transcripts (four from heuristic group and four from vocabulary group). After discussing and incorporating codes, we created a codebook that the two researchers agreed on. Using this codebook, the researchers divided the remaining transcripts, each coding eight transcripts per condition. Like many other qualitative analyses in S&P research [71, 72], the two researchers discussed and resolved coding conflicts in several weekly meetings. In this analysis, as we prioritized identifying emerging themes, we did not calculate the inter-rater reliability (IRR) to seek theoretical agreement [73].
@@ 5.2.3 Quantitative Results
Improved Sketching Efficiency. As shown in Table VI, heuristic-guided participants reached their satisfied designs faster (average of 11.53 v.s. 14.80 minutes) and covered more privacy-related design decisions (average of 11.5 v.s. 8.9 per sketch), compared to the vocabulary group.
@@ 5.3.1 Data Collection
As shown in Figure 6, each participant completed two interpretation tasks, each with a 7-minute time limit. In each task, they received a randomized list of privacy-related design decisions (excluding the “Procedure” column in Table XI) and an unannotated copy of a sketch created by a prior participant. We then asked them to annotate and orally describe the privacy-related design decisions they identified. This annotation is similar to what they have done during the creation tasks (Section 5.2.1). Table 5.2.4 outlines the interpretation task order, where each participant interpreted two sketches of different scenarios: one created by a heuristic-guided participant and another from the vocabulary group. We maintained scenario order consistent with the creation tasks (Table 5.2.1) to avoid biases from the most recent scenario they had sketched. For example, after sketching for Afterlife Chatbot as his second creation, P3 would first interpret a sketch of another scenario. After each task, we recorded their interpretation time and asked them to complete a NASA TLX survey [69].
@@ 5.3.3 Quantitative Results
Eased Workload of Interpretation. All participants finished their interpretation within the time limit. There was no significant difference in interpretation time (4.9 minutes for heuristic-guided sketches v.s. 5.1 minutes for vocabulary-guided sketches). However, participants consistently reported significantly lower mental, physical, temporal, and effort-related demands and reduced frustration when working with heuristic-guided sketches. They also perceived improved performance, as indicated by Wilcoxon Signed-Rank tests (Table X). Furthermore, participants identified more privacy design decisions in heuristic-guided sketches, with an average of 9.4 elements per sketch compared to 8.1 from vocabulary-guided ones. These results suggest that heuristic-guided sketches are more straightforward to interpret, decreasing workload while enhancing outcomes.
@@
At the level of a complete privacy design, heuristic-guided sketches demonstrated higher per-sketch interpretation performance, again with a higher precision (79.1% v.s. 44.6% for the vocabulary group), reflecting the proportion of matches among all interpretations in a sketch, and higher recall (71.1% v.s. 40.5%), representing the percentage of alignments among the creator’s all design decisions. These differences are all statistically significant according to the Wilcoxon Signed-Rank tests [74], highlighting the improved communication efficiency facilitated by heuristic guidance.
@@ 6 Discussion & Future Work
Sketches v.s. Diagrams. We experimented with both sketches and diagrams in this project. We found that complex diagram structures hindered the learning process (Section 4.1.4). For instance, when students used diagrams to illustrate interactions between stakeholders and data flows, they needed a thorough understanding of diagram vocabulary. This steep learning curve increased their cognitive load, making it harder to complete the privacy design.
@@
Sketches and Threat Modeling. Our current study focuses on facilitating the communication of privacy design concepts through sketches. Future research may further explore the benefits of sketches in other security and privacy applications. For example, participants may develop a stronger awareness of privacy risks during sketching [81, 82, 83, 84]. The improved communication enabled by privacy sketches may enhance the collaborative decision-making process [1]. Conventional threat modeling approaches [85, 86] often rely on static tables or data flow diagrams (DFDs) to define technical scope and decompose applications. Compared to tables and DFDs, sketching may better capture dynamic aspects of a system, such as stakeholder interactions and contextual nuances that are difficult to convey textually.
@@ 7 Limitations
Ground Truth for the Creation Task. The creation task has no ground truth. While our codebook (Table XI) provided a structured approach for labeling and interpreting data, it may have constrained the granularity of our analysis. By relying on predefined categories, we may have overlooked subtle contextual differences or nuances in participant responses.
@@ Acknowledgement
We thank the anonymous reviewers and the shepherd for their invaluable feedback and the participants for their kind involvement in our studies.
@@ F Neither responded
True Positive (TP) Interpreter’s response fully or partially aligned with the creator’s response (partial alignment weighted by $\epsilon$). #$\#A+\epsilon\cdot\#B$
False Positive (FP) Interpreter responded when the creator did not, including the unaligned ($1-\epsilon$) part of $B$. #$\#D+(1-\epsilon)\cdot\#B$
False Negative (FN) The creator responded, but the interpreter either did not respond or misinterpreted the response. #$\#C+\#E$
True Negative (TN) Neither the creator nor the interpreter responded. #$\#F$
Precision The proportion of correctly aligned responses out of all responses made by the interpreter. #$\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#D+(1-\epsilon)\cdot\#B}$
Recall The proportion of correctly aligned responses out of all responses made by the creator. #$\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#C+\#E}$

### v2

@@ Abstract
00footnotetext:  00footnotetext: Equal contributions.00footnotetext: This report is an extended version of an IEEE S&P 2025 paper [References]. We include Table XIV to further prove the improved communication efficiency of heuristic-guided privacy design sketches (Section 5.3.3). This table was skipped in the original version due to space constraints.
@@ 1 Introduction
We conducted a between-subjects experiment to validate the effectiveness of the heuristic-based approach (Section 5). We recruited a total of 24 participants, with each participant creating a few sketches and interpreting the sketches from other participants. We found that participants who received heuristic-based instructions were able to cover 30% more of the privacy-related decisions in their sketches than the participants without exposure to these instructions. The final sketches from the participants who received heuristic-based instruction are also more readable, with an increase of 77% in interpretation accuracy.
@@ 2.2 Education for Privacy Design
Current privacy design education primarily focuses on privacy concepts [16] and design methodologies [17, 18, 19], with most educators using case studies [17, 20] as the main form of instruction. For instance, some classes include a privacy review procedure [21, 20], where students assess smartphone privacy notices, identify issues, and propose design improvements in brief essays. Organizations such as the National Institute of Standards and Technology (NIST) [22] and the International Association of Privacy Professionals (IAPP) [23] also provide privacy design education by offering standards, certifications, and online learning resources, which provide practical guidance and structure for both academic and professional learning. Additionally, educators have expanded the instruction of privacy design to a broader audience, including industry professionals and policymakers [24]. In contrast, our work aims to complement existing curricula by exploring methods to teach students how to sketch privacy designs, equipping them with privacy design communication skills.
@@ 2.3 Diagrams for Privacy Design
Practitioners often use a combination of diagrams to communicate privacy considerations in system designs [3, 4, 44, 45]. Research has shown that these diagrams help developers visualize system architecture by simplifying complex design elements, making it easier to communicate and design privacy-aware applications [4]. For example, when tasked with communicating the privacy design of an IoT application for diabetes treatment and monitoring, developers used Data Flow Diagrams (DFDs) [46] and Unified Modeling Language (UML) [47] to break down the flow of sensitive health data into manageable components, with annotations clarifying the encryption mechanism [4].
While existing studies shed light on the use of diagrams for communicating privacy considerations in system design, their focus has been on professional developers. How to teach these diagramming techniques to novice data science students remains unclear. Our work aims to address this gap by using sketches as the lens to investigate the specific challenges students face when communicating privacy design through diagrams and provide insights on targeted guidance to address these difficulties.
@@ 2.4 Developer Support for Privacy Design
Prior research has explored many approaches to support experienced developers in privacy design [48, 49]. For instance, some studies have offered insights into how to help developers make privacy design choices by identifying the common challenges in understanding and implementing privacy considerations [50, 51, 52]. Additionally, various tools have been developed to support developers by embedding privacy practices directly into development workflows. For example, Coconut, an Android Studio plugin, helps developers create privacy-friendly apps by requiring privacy annotations and assisting them in organizing privacy-related details [53]. Similarly, PARROT helps developers create privacy-aware IoT applications by providing interactive privacy annotations and guidance [4].
@@ 4.1.1 Design
Second, we offered preset attribute options to help novice students explore the design space systematically (Figure 3 and 4). Instead of describing each DataAction in natural language, students could select predefined attribute options. For example, when specifying data source to Collect, students could choose “direct” for newly generated data or “secondary” for pre-existing data.
@@ 4.3.3 Observations
All participants completed the warm-up within 20 minutes (Mean = 18.5 minutes, SD = 1.2 minutes) and reached satisfaction with each design within 15 minutes (Mean = 13.6 minutes, SD = 2.9 minutes). First, we observed a significant reduction in completion time with the vocabulary-based approach (14.3 vs. 18.6 minutes in Section 4.2.4), as participants reported that the warm-up session helped them become familiar with the method. Additionally, there was a clear preference for heuristics over vocabulary-based design. For example, participants who started with heuristic-guided sketching often intentionally applied the multi-layered heuristic when sketching vocabulary-based designs. However, we did not observe a tendency to re-apply visual vocabulary to heuristic-based tasks. T13 noted, “I felt freer to move beyond ellipses and boxes and to organize my ideas in a more trackable way.”
@@ 5.1 Study Overview
Study Procedure and Apparatus. We followed a between-subjects study design. Participants received either vocabulary-based or heuristic-based guidance for their tasks. They completed tasks on an iPad with a stylus using a sketch board application (e.g., Notability [62]).
@@
Then, each participant was asked to interpret two sketches created by previous participants (Table 5.3). Finally, we interviewed participants about what hindered or assisted them in their creation and interpretation tasks.
@@
Annotating Visual Sketches. Comparing high-dimensional sketches is challenging. To address this, we developed a codebook (Appendix Table LABEL:tab:design_hints) to annotate privacy-related design decisions covered in the sketches. We then used the annotated decisions to assess the design coverage in sketch creation and the communication effectiveness between the sketch creators and interpreters (Sections 5.2, 5.3).
@@ 5.2.1 Data Collection
We presented participants with a list of privacy-related design decisions (Table LABEL:tab:design_hints) and asked them to describe each decision they had included, annotating the corresponding content on their sketches. We didn’t set time limits for this process. Participants could skip any decisions they felt were not covered but could not modify their original designs.
@@ 5.2.2 Data Analysis
There is no perfect sketch for the creation task, as participants may have different designs for the same scenario, which may evolve during sketching. So, we use the codebook (Table LABEL:tab:design_hints) to systematically capture the unique design decisions reflected in each sketch. To assess the quality of the sketches, we documented the privacy design decisions annotated by participants, counted the total number of decisions included in each sketch, and calculated the coverage percentage of each design decision across all sketches in each condition.
Next, we conducted a thematic analysis [73] of post-task interview transcripts and notes we took during the study. First, two researchers independently coded eight transcripts (four from the heuristic group and four from the vocabulary group). After discussing and incorporating codes, we created a codebook that the two researchers agreed on. Using this codebook, the researchers divided the remaining transcripts, each coding eight transcripts per condition. Like many other qualitative analyses in S&P research [74, 75], the two researchers discussed and resolved coding conflicts in several weekly meetings. In this analysis, as we prioritized identifying emerging themes, we did not calculate the inter-rater reliability (IRR) to seek theoretical agreement [76].
@@ 5.2.3 Quantitative Results
Improved Sketching Efficiency. As shown in Table VI, heuristic-guided participants reached their satisfied designs faster (average of 11.53 vs. 14.80 minutes) and covered more privacy-related design decisions (average of 11.5 vs. 8.9 per sketch), compared to the vocabulary group.
@@ 5.3.1 Data Collection
As shown in Figure 6, each participant completed two interpretation tasks, each with a 7-minute time limit. In each task, they received a randomized list of privacy-related design decisions (excluding the “Procedure” column in Table LABEL:tab:design_hints) and an unannotated copy of a sketch created by a prior participant. We then asked them to annotate and orally describe the privacy-related design decisions they identified. This annotation is similar to what they have done during the creation tasks (Section 5.2.1). Table 5.3 outlines the interpretation task order, where each participant interpreted two sketches of different scenarios: one created by a heuristic-guided participant and another from the vocabulary group. We maintained scenario order consistent with the creation tasks (Table 5.2.1) to avoid biases from the most recent scenario they had sketched. For example, after sketching for Afterlife Chatbot as his second creation, P3 would first interpret a sketch of another scenario. After each task, we recorded their interpretation time and asked them to complete a NASA TLX survey [72].
@@ 5.3.3 Quantitative Results
Eased Workload of Interpretation. All participants finished their interpretation within the time limit. There was no significant difference in interpretation time (4.9 minutes for heuristic-guided sketches vs. 5.1 minutes for vocabulary-guided sketches). However, participants consistently reported significantly lower mental, physical, temporal, and effort-related demands and reduced frustration when working with heuristic-guided sketches. They also perceived improved performance, as indicated by Wilcoxon Signed-Rank tests (Table X). Furthermore, participants identified more privacy design decisions in heuristic-guided sketches, with an average of 9.4 elements per sketch compared to 8.1 from vocabulary-guided ones. These results suggest that heuristic-guided sketches are more straightforward to interpret, decreasing workload while enhancing outcomes.
@@
At the level of a complete privacy design, heuristic-guided sketches demonstrated higher per-sketch interpretation performance, again with a higher precision (79.1% vs. 44.6% for the vocabulary group), reflecting the proportion of matches among all interpretations in a sketch, and higher recall (71.1% vs. 40.5%), representing the percentage of alignments among the creator’s all design decisions. These differences are all statistically significant according to the Wilcoxon Signed-Rank tests [77], highlighting the improved communication efficiency facilitated by heuristic guidance.
@@ 6 Discussion & Future Work
Sketches vs. Diagrams. We experimented with both sketches and diagrams in this project. We found that complex diagram structures hindered the learning process (Section 4.1.4). For instance, when students used diagrams to illustrate interactions between stakeholders and data flows, they needed a thorough understanding of diagram vocabulary. This steep learning curve increased their cognitive load, making it harder to complete the privacy design.
@@
Sketches and Threat Modeling. Our current study focuses on facilitating the communication of privacy design concepts through sketches. Future research may explore the benefits of sketches in other security and privacy applications. For example, participants may develop a stronger awareness of privacy risks during sketching [84, 85, 86, 87]. The improved communication enabled by privacy sketches may enhance the collaborative decision-making process [1]. Conventional threat modeling approaches [88, 89] often rely on static tables or Data Flow Diagrams (DFDs) to define technical scope and decompose applications. Compared to tables and DFDs, sketching may better capture dynamic aspects of a system, such as stakeholder interactions and contextual nuances that are difficult to convey textually.
@@ 7 Limitations
Ground Truth for the Creation Task. The creation task has no ground truth. While our codebook (Table LABEL:tab:design_hints) provided a structured approach for labeling and interpreting data, it may have constrained the granularity of our analysis. By relying on predefined categories, we may have overlooked subtle contextual differences or nuances in participant responses.
@@ Acknowledgement
We thank the anonymous reviewers and the shepherd for their invaluable feedback and the participants for their kind involvement in our studies. This research is in part supported by a gift from Solix Technologies, the National Science Foundation CNS-2341187, CNS-2426397, a Google PSS Faculty Award, and a Meta Research Award.
@@ F Neither responded
True Positive (TP) Interpreter’s response fully or partially aligned with the creator’s response (partial alignment weighted by $\epsilon$). $A+\epsilon\cdot\#B$
False Positive (FP) Interpreter responded when the creator did not, including the unaligned ($1-\epsilon$) part of $B$. $D+(1-\epsilon)\cdot\#B$
False Negative (FN) The creator responded, but the interpreter either did not respond or misinterpreted the response. #C+#E\#C+\#E
True Negative (TN) Neither the creator nor the interpreter responded. #F\#F
Precision The proportion of correctly aligned responses out of all responses made by the interpreter. $\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#D+(1-\epsilon)\cdot\#B}$
Recall The proportion of correctly aligned responses out of all responses made by the creator. $\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#C+\#E}$
@@
Appendix D Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
D.1 Summary
The paper presents a user study of n=24 data science students (both undergraduate and graduate) who were asked to sketch privacy designs of a set of privacy scenarios after being given three heuristics to follow at the beginning of the study. The paper describes two pre-studies. The first pre-study was a need-finding study (n=12) that found that students lacked the vocabulary needed to sketch privacy designs effectively, the sketching may require multiple iterations, and the students had difficulty planning the sketch space in advance. The second pre-study is a teaching methods experiment that investigates 3 methods: object-oriented diagramming, vocabulary-based sketching, and heuristic-based sketching. The findings suggested that heuristic-based and vocabulary-based sketching would be the best for data science students, and those methods are examined in the main study. The main study (n=24) compared the two methods in a sketching task, followed by a post-task interview. The findings suggest that the simple heuristic method performed best in sketching efficiency and broader coverage of privacy-related design decisions and had a reduced learning curve. A heuristic-based approach combined with the multi-layer representation is recommended for privacy design sketches conducted by data science students.
D.2 Scientific Contributions
  • Provides a Valuable Step Forward in an Established Field
  • Establishes a New Research Direction
D.3 Reasons for Acceptance
  1. 1.
    Provides a Valuable Step Forward in an Established Field. The paper contributes to the research on sketching privacy designs and offers insights into how we should teach data science students about data privacy design.
  2. 2.
    Establishes a New Research Direction. This paper presents the design and evaluation of a novel approach to prompting data science students to consider privacy in their design. The proposed heuristic-based prompting offers a novel direction of investigation for future work.